# Automated discovery and optimization of autonomous quantum error correction codes for a general open quantum system

### v1

@@ III.3 Ququart with photon loss
We now consider the lowest four energy levels of a harmonic oscillator (which can be treated as a ququart) with photon-loss decay, i.e. a decay jump operator given by the appropriately truncated harmonic oscillator annihilation operator
@@
We assume that there are no other decoherence channels. We ran the optimal AQEC code search algorithm on this case. The calculations consistently converged to the code space $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$, with $\hat{b}=\hat{O}=0$. In other words, the optimal code found by the algorithm is to do nothing.
We then ask what happens if we apply the 13 code of Sec. III.B in this case. Even when the decay jump operator is given by Eq. (18), the states $\left|1\right\rangle$ and $\left|3\right\rangle$ are stabilized effectively by the operator $\hat{b}$ in Eq. (17), which quickly brings back any population that decays into the error states $\left|0\right\rangle$ and $\left|2\right\rangle$, respectively. The key matrix elements to analyze are then $\rho_{02}$ and $\rho_{13}$, or equivalently $\rho_{20}$ and $\rho_{31}$. The equations governing the dynamics of these matrix elements are
$\dot{\rho}_{02}$ $=$ $\sqrt{3}\gamma\rho_{13}-\Gamma\rho_{02}-\frac{\gamma}{\sqrt{2}}\rho_{02}$
@@
The last term in the equation for $\rho_{02}$ arises from the relaxation of the state $\left|2\right\rangle$ into the state $\left|1\right\rangle$, which is an uncorrectable error that we expect to make only a small contribution to the effective decoherence dynamics. We therefore ignore this term. The terms with the coefficient $\Gamma$ correspond to the fast reexcitation of the error states $\left|0\right\rangle$ and $\left|2\right\rangle$ back to the code words $\left|1\right\rangle$ and $\left|3\right\rangle$, respectively. We are then left with the following situation: $\rho_{13}$ decays at the rate $2\gamma$ and is replenished at the rate $\sqrt{3}\gamma$ that appears in the first term in the equation for $\rho_{02}$. The net effect is that $\rho_{13}$ decays at the rate $(2-\sqrt{3})\gamma$. In other words, the effect of applying this AQEC code is to replace the free-evolution situation in which we have only relaxation with rate $\gamma$ by an effective situation in which we have mainly pure dephasing between the code words with a rate of roughly $\gamma/4$. Using Eq. (15), the fidelity in this situation is expected to be given by $F\approx 0.88$, which is significantly higher than the fidelity $F_{0}\approx 0.65$. It is then surprising that the algorithm did not find the active AQEC code. This result suggests that there are local optima in the fidelity landscape.
Figure 5: Code fidelity ($F$) and decay rate suppression factor $\kappa$ for the optimal AQEC code for a 4-level system with decay jump operator given by Eq. (20). The case $\alpha=0$ corresponds to uniform decay rates, where the 13 code is expected to be optimal. The case $\alpha=0.5$ corresponds to photon loss decay. Starting the search from a random guess (random seed) did not produce any nontrivial codes. Interestingly, nontrivial codes appear as soon as we move away from $\alpha=0.5$. Furthermore, using the 13 code as a starting point for the search produces a nontrivial code even at $\alpha=0.5$.
To further investigate the difference between the uniform decay case treated in Sec. III.B and the photon loss case treated in this subsection, we ask what happens in intermediate cases. In particular, we consider the jump operator
where the exponent $\alpha$ is between 0 (uniform decay case) and 0.5 (photon loss case). The fidelity as a function of $\alpha$ is plotted in Fig. 5. When starting with random seeds (i.e. random guesses) for the search, we ran the algorithm 5 times for each setting and chose the highest fidelity, noting here that most runs produced essentially the same results. Interestingly, even for $\alpha=0.49$, the algorithm consistently finds a nontrivial AQEC code with fidelity $F\approx 0.8867$, which is close to the value mentioned above when applying the 13 code. Indeed, if we inspect the codes found by the algorithm for all values of $\alpha$ excluding 0.5, we find that the code space is approximately $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$. If we start the search from the 13 code, we find that the fidelity agrees with that found by the random-seed search for most values of $\alpha$. One exception is the case $\alpha=0.5$, where the random-seed search fails to find the optimal code. Another exception is the case of small $\alpha$, where starting the search from a very good guess gives a higher fidelity than the random seed search for the same number of optimization iterations.
III.4 Photon loss in higher-dimensional qudit
@@
Figure 6: Progression of AQEC code for 5-level system with photon-loss decay. This figure shows an instance in which the algorithm failed to produce a nontrivial AQEC code. The fidelity converges to the free-evolution value $F_{0}$ (shown as the horizontal dotted line in Panel a). Panel (b) shows the probability for the code to be outside two potential code spaces, namely the space of the binomial code and the space $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$. The code space gradually approaches $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$, which is a clear indication that the algorithm is converging towards the trivial code of effectively doing nothing. This figure is representative of most runs of the algorithm. Figure 7 shows one of the relatively rare instances where the algorithm found a nontrivial code.
@@
Figure 7: Progression of AQEC code for 5-level system with photon-loss decay. This figure shows an instance in which the algorithm found an AQEC code with $F>F_{0}$. Such successful instances occurred in a few percent of the runs, each starting from a random initial guess. Although the fidelity is $F=0.998$ for the code reached after $10^{5}$ iterations, the code space is significantly different from that of the binomial code. Panel (a) suggests that there is still plenty of room for further optimizing the code. On the other hand, the high fidelity indicates that this code is already a very good AQEC code. In some sense, this instance represents a case in which the algorithm discovered a new code that is very good, although it is still not as good as the binomial code, which has $1-F=6\times 10^{-6}$. As can be seen in Panel (c), this code requires that a large number of transitions are induced with varying transition rates.
@@
Next we consider the lowest five energy levels of a harmonic oscillator with photon-loss decay, i.e. the decay jump operator being given by the truncated annihilation operator
@@
This system allows the implementation of the bosonic binomial code. The code space is spanned by the states $(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2}$ and $\left|2\right\rangle$. The annihilation operator takes these states into the states $\left|3\right\rangle$ and $\left|1\right\rangle$, respectively, both with the rate $2\gamma$.
@@
[…]
@@
where $\delta\theta=\pi/2-2\arctan\left(e^{-\gamma\delta t}\right)=\gamma\delta t+O(\delta t^{2})$ and $\hat{\sigma}_{y}^{(04)}=i\left(\left|4\right\rangle\left\langle 0\right|-\left|0\right\rangle\left\langle 4\right|\right)$. This operation can in turn be implemented by adding to the Hamiltonian the time-independent operator $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$. In other words, $O_{0,4}=-O_{4,0}=-i$, with all other matrix elements in $\hat{O}$ equal to zero.
If we take the code space $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and the matrices $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we set $\gamma\tau=1$, we obtain the fidelity $F=0.999994=1-6\gamma/\Gamma$. If we run the optimization algorithm using the above settings, the fidelity increases by less than $\sim 10^{-9}$ in $10^{3}$ iterations, at which point the algorithm meets our early termination conditions.
If we fix $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we start the search from a randomly generated basis, the fidelity comes within $10^{-8}$ of the maximum value after about 10-15 iterations. In other words, as in the previous subsections, the code basis converges quickly. Similarly, if we fix the code space to $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we optimize $\hat{O}$, the fidelity quickly rises and saturates within about 10-15 iterations. The maximum fidelity is typically $10^{-6}$ lower than the maximum value of $F=0.999994$. As with the code basis, the Hamiltonian operator $\hat{O}$ converges quickly. If we fix the code space to $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$, and we optimize $\hat{b}$, the fidelity reaches $F=0.9998$ after $10^{5}$ iterations. When compared with the other two components of the AQEC code, the induced decay operator $\hat{b}$ is the slowest converging component.
@@
IV Five-level system with perturbed-photon-loss decay law
Figure 8: An example of optimizing an AQEC code for a slightly perturbed system. The infidelity $(1-F)$ and the decay rate suppression factor $\kappa$ are plotted as functions of the exponent $\alpha$ in Eq. (24). When $\alpha=0.5$ (unperturbed case), the binomial code performs better than any AQEC code that we obtained numerically starting from a random initial guess. In this figure we took the best result out of ten runs using random initial guesses. When $\alpha=0.45$, applying the binomial code in its original form gives a relatively low fidelity. Starting from the binomial code and numerically optimizing it gives a better fidelity than any code obtained from a random initial guess. When $\alpha=0.4$ the best AQEC obtained from a random initial guess performs better than the code obtained by optimizing the binomial code, indicating that the perturbation is so large that the binomial code is no longer a useful starting point for optimization.
We now consider a situation similar to that treated in Sec. III.D, but with system parameters that are not slightly different from those of an ideal harmonic oscillator. Indeed, in a general qudit or anharmonic oscillator, we do not expect the parameters to follow perfectly regular patterns. In such situations, intuition can at best produce an approximate error correction protocol. Numerical optimization is then particularly useful to optimize the protocol and achieve the best possible performance.
@@ IV Five-level system with perturbed-photon-loss decay law
$\hat{a}=\sqrt{\gamma}\left(\begin{array}[]{ccccc}0&1&0&0&0\\ 0&0&2^{\alpha}&0&0\\ 0&0&0&3^{\alpha}&0\\ 0&0&0&0&4^{\alpha}\end{array}\right),$ (24)
@@
with $\alpha=0.45$ or $\alpha=0.4$. The results are shown in Fig. 8. If we take the binomial code in its original form and apply it to the cases $\alpha=0.45$ and $\alpha=0.4$, we obtain the fidelity values $F=0.9967$ and $F=0.988$, respectively. While these are high numbers, the fidelity can be improved further with numerical optimization. We therefore take the ideal binomial code and use it as the starting point for numerical optimization. We now obtain the fidelity values $F=0.99957$ and $0.9983$, respectively. In both cases, the optimization leads to a reduction in the effective decay rate by a factor of about 9. In the case $\alpha=0.45$, starting the optimization from the binomial code produced the highest fidelity and lowest effective decoherence rate among all our calculations. In the case $\alpha=0.4$, the best code that we found was obtained from a random-seed search, indicating that the perturbation is so large that the binomial code does not provide a computational advantage as a starting point for the search.
V Discussion

### v2

@@ IV.3 Ququart with photon loss
We now consider the lowest four energy levels of a harmonic oscillator (which can be treated as a ququart) with photon loss decay, i.e. a decay jump operator given by the appropriately truncated harmonic oscillator annihilation operator
@@
We assume that there are no other decoherence channels. We ran the optimal AQEC code search algorithm on this case. When we initialized the AQEC code components to random values, the calculations consistently converged to the code space $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$, with $\hat{b}=\hat{O}=0$. In other words, the optimal protocol found by the algorithm is the one where we do nothing.
We then ask what happens if we apply the 13 code of Sec. IV.B in this case. Even when the decay jump operator is given by Eq. (18), the states $\left|1\right\rangle$ and $\left|3\right\rangle$ are stabilized effectively by the operator $\hat{b}$ in Eq. (17), which quickly brings back any population that decays into the error states $\left|0\right\rangle$ and $\left|2\right\rangle$, respectively. The key matrix elements to analyze are then $\rho_{02}$ and $\rho_{13}$, or equivalently $\rho_{20}$ and $\rho_{31}$. The equations governing the dynamics of these matrix elements are
$\dot{\rho}_{02}$ $=$ $\sqrt{3}\gamma\rho_{13}-\Gamma\rho_{02}-\gamma\rho_{02}$
@@
The last term in the equation for $\rho_{02}$ arises from the relaxation of the state $\left|2\right\rangle$ into the state $\left|1\right\rangle$, which is an uncorrectable error that will make only a small contribution to the effective decoherence dynamics when $\Gamma\gg\gamma$. We therefore ignore this term. The terms with the coefficient $\Gamma$ correspond to the fast reexcitation of the error states $\left|0\right\rangle$ and $\left|2\right\rangle$ back to the code words $\left|1\right\rangle$ and $\left|3\right\rangle$, respectively. We are then left with the following situation: $\rho_{13}$ decays at the rate $2\gamma$ and is replenished at the rate $\sqrt{3}\gamma$ that appears in the first term in the equation for $\rho_{02}$. The net effect is that $\rho_{13}$ decays at the rate $(2-\sqrt{3})\gamma$. In other words, the effect of applying this AQEC code is to replace the free-evolution situation in which we have only relaxation with rate $\gamma$ by an effective situation in which we have mainly pure dephasing between the code words with a rate of roughly $\gamma/4$. Using Eq. (15), the fidelity in this situation is expected to be given by $F\approx 0.88$, which is significantly higher than the fidelity $F_{0}\approx 0.65$. It is then surprising that the algorithm did not find the active AQEC code. This result suggests that there are local optima in the fidelity landscape.
It is worth noting here that the relatively modest improvement described in the previous paragraph, from relaxation with rate $\gamma$ to pure dephasing with rate $\gamma/4$, is unlikely to materialize in a realistic experiment, considering that the driving protocol in any realistic setup will contribute its own noise and, partially or fully, negate the improvement gained by the error correction mechanism.
It is also interesting to consider the generalization of the 13 code dynamics described above, as well as the code found in Ref. Zeng , to the scenario where the code space is composed of the Fock states $\left|N\right\rangle$ and $\left|N+2\right\rangle$ (or more generally $\left|N_{1}\right\rangle$ and $\left|N_{2}\right\rangle$ with $\left|N_{1}-N_{2}\right|\geq 2$) of a harmonic oscillator. The equations of motion for the relevant off-diagonal matrix elements become
$\dot{\rho}_{N-1,N+1}$ $=$ $\sqrt{N(N+2)}\gamma\rho_{N,N+2}-\Gamma\rho_{N-1,N+1}-N\gamma\rho_{N-1,N+1}$
$\dot{\rho}_{N,N+2}$ $=$ $-(N+1)\gamma\rho_{N,N+2}+\Gamma\rho_{N-1,N+1}.$ (20)
Following the same argument as above, we find that the effective dynamics is that of pure dephasing in which $\rho_{N,N+2}$ decays at the rate $\left(N+1-\sqrt{(N+1)^{2}-1}\right)\gamma$, which can be approximated as $\gamma/\left(2\sqrt{N}\right)$ for large $N$. One can therefore, in theory, reduce the effective decay rate to arbitrarily small values by using very large photon numbers, keeping in mind that the induced decay rate $\Gamma$ must now satisfy the condition $\Gamma\gg N\gamma$. The reason behind the suppression of the effective decay rate is that, in the limit $N\rightarrow\infty$, the decay rates of the states $\left|N\right\rangle$ and $\left|N+2\right\rangle$ become essentially equal, which leads to the simple scenario described in Sec. IV.B.
IV.4 Photon loss in higher-dimensional qudit
@@
Figure 3: Progression of AQEC code with optimization iteration number $k$ for 5-level system with photon loss decay. This figure shows an instance in which the algorithm failed to produce a nontrivial AQEC code. The fidelity converges to the free-evolution value $F_{0}$ (shown as the horizontal dotted line in Panel a). Panel (b) shows the probability for the code to be outside two code spaces that can be intuitively guessed as candidate code spaces, namely the space of the binomial code and the space $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$. The code space gradually approaches $\left\{\left|0\right\rangle,\left|1\right\rangle\right\}$. All the plots in this figure indicate that the algorithm is converging towards the trivial code of effectively doing nothing. This figure is representative of most runs of the algorithm. Figure 4 shows one of the relatively rare instances where the algorithm found a nontrivial code.
@@
Figure 4: Same as Fig. 3, but for an instance in which the algorithm succeeded in finding an AQEC code with $F>F_{0}$. Such successful instances occurred in a few percent of the runs, each starting from a random seed. Although the fidelity is quite high ($F=0.998$) after $10^{5}$ iterations, Panel (b) clearly shows that the code is significantly different from the binomial code, which we expect is the optimal code in this model. Panel (a) suggests that there is still plenty of room to further optimize the code. On the other hand, the high fidelity indicates that this code is already a very good AQEC code. In some sense, this instance represents a case in which the algorithm discovered a new code that is very good, although it is still not as good as the binomial code, which has $1-F=6\times 10^{-6}$. As can be seen in Panel (c), this code requires that a large number of transitions are induced with varying transition rates. The legend specifies only the highest eight lines.
@@
Next we consider the lowest five energy levels of a harmonic oscillator with photon loss decay, i.e. the decay jump operator being given by the truncated annihilation operator
@@
This system allows the implementation of the bosonic binomial code Michael . The code space is spanned by the states $(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2}$ and $\left|2\right\rangle$. The annihilation operator takes these states into the states $\left|3\right\rangle$ and $\left|1\right\rangle$, respectively, both with rate $2\gamma$.
@@
This operator does indeed return the state after an error to the original state. However, one more error-correction component is needed. When no photon loss occurs during a time $\delta t$, the state $(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2}$ is deformed and becomes $c(\left|0\right\rangle+e^{-\gamma\delta t}\left|4\right\rangle)$, where $c$ is a normalization constant. To return this state to an equal superposition of $\left|0\right\rangle$ and $\left|4\right\rangle)$, we must apply a correction operation. This correction can be achieved via the unitary operator
@@
where $\delta\theta=\pi/2-2\arctan\left(e^{-\gamma\delta t}\right)=\gamma\delta t+O(\delta t^{2})$ and $\hat{\sigma}_{y}^{(04)}=i\left(\left|4\right\rangle\left\langle 0\right|-\left|0\right\rangle\left\langle 4\right|\right)$. This operation can in turn be implemented by adding to the Hamiltonian the time-independent operator $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$. In other words, $O_{0,4}=-O_{4,0}=-i$, with all other matrix elements in $\hat{O}$ equal to zero. It should be noted that the operator $\hat{\sigma}_{y}^{(04)}$ involves a four-photon process and therefore seems challenging to realize experimentally. There are, however, theoretical proposals and initial experimental results on implementing such multi-photon operations Kwon2022 ; Chang ; Eriksson .
If we take the code space $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and the matrices $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we set $\gamma\tau=1$, we obtain the fidelity $F=0.999994=1-6\gamma/\Gamma$. If we run the optimization algorithm using the above settings, the fidelity increases by less than $\sim 10^{-9}$ in $10^{3}$ iterations, at which point the algorithm meets our early termination condition.
If we fix $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we start the search from a randomly generated basis, the fidelity comes within $10^{-8}$ of the maximum value after about 10-15 iterations. In other words, the code basis converges quickly. Similarly, if we fix the code space to $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and $\hat{b}$ as in Eq. (22) with $\Gamma/\gamma=10^{6}$, and we optimize $\hat{O}$, the fidelity quickly rises and saturates within about 10-15 iterations. The maximum fidelity is typically $10^{-6}$ lower than the maximum value of $F=0.999994$. As with the code basis, the Hamiltonian operator $\hat{O}$ converges quickly. If we fix the code space to $\left\{(\left|0\right\rangle+\left|4\right\rangle)/\sqrt{2},\left|2\right\rangle\right\}$ and $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$, and we optimize $\hat{b}$, the fidelity reaches $F=0.9998$ after $10^{5}$ iterations. When compared with the other two components of the AQEC code, the induced decay operator $\hat{b}$ is the slowest-converging component. In other words, the convergence behaviour of the individual AQEC code components is similar to that observed in the case of the ququart.
@@
V Five-level system with perturbed-photon-loss decay law
Figure 5: Optimizing AQEC code for a slightly perturbed system. The infidelity (1F)(1-F) and the decay rate suppression factor $\kappa$ are plotted as functions of the exponent $\alpha$ in Eq. (24). The data labeled “binomial” are obtained by applying the binomial code in its original form to the perturbed system. The data labeled “optimized” are obtained by optimizing the AQEC code starting from the binomial code as an initial guess. For the data labeled “random seed”, we took the best result out of ten runs using random seeds for each value of $\alpha$. When $\alpha=0.5$ (unperturbed case), the binomial code performs better than any AQEC code that we obtained numerically starting from a random seed. Furthermore, numerical optimization did not improve on the binomial code. When $\alpha=0.45$, applying the binomial code in its original form gives a relatively low fidelity. Starting from the binomial code and numerically optimizing it gives a better fidelity than any code obtained from a random seed. When $\alpha=0.4$ the best AQEC obtained from a random seed performs better than the code obtained by optimizing the binomial code, indicating that the perturbation is so large that the binomial code is no longer a useful starting point for optimization. These results demonstrate how our algorithm can find optimal AQEC codes that cannot be obtained by simple intuition about the system.
We now consider a situation similar to that treated in Sec. IV.D, but with system parameters that are slightly different from those of an ideal harmonic oscillator. In a realistic setup, we do not expect the parameters to follow perfectly regular patterns. For example, in a superconducting circuit, there will inevitably by nonlinear terms that shift the Hamiltonian away from the perfect harmonic oscillator form. In such situations, intuition can at best produce an approximate error correction protocol. Numerical optimization is then particularly useful to optimize the protocol and achieve the best possible performance. It should be noted here that the perturbation can typically be characterized experimentally, such that we can know precisely what perturbation we are dealing with. In other words, we consider only static perturbations, not dynamic noise that is unpredictable in every run of the experiment.
@@ V Five-level system with perturbed-photon-loss decay law
$\hat{a}=\sqrt{\gamma}\left(\begin{array}[]{ccccc}0&1&0&0&0\\ 0&0&2^{\alpha}&0&0\\ 0&0&0&3^{\alpha}&0\\ 0&0&0&0&4^{\alpha}\\ 0&0&0&0&0\end{array}\right),$ (24)
@@
with $\alpha$ slightly below 0.5. The results are shown in Fig. 5. If we take the binomial code in its original form and apply it to the cases $\alpha=0.45$ and $\alpha=0.4$, we obtain the fidelity values $F=0.9967$ and $F=0.988$, respectively. While these are high numbers, indicating that the binomial code is still a good AQEC code in the presence of the perturbation, the fidelity can be improved further with numerical optimization. We therefore take the original binomial code and use it as the initial guess for the automated code optimization algorithm. We now obtain the fidelity values $F=0.99957$ and $0.9983$, respectively. In both cases, the optimization leads to a reduction in the effective decay rate by a factor of about 9. In the case $\alpha=0.45$, starting the optimization from the binomial code produced the highest fidelity and lowest effective decoherence rate among all our calculations. In the case $\alpha=0.4$, the best code that we found was obtained from a random-seed search, indicating that the perturbation is so large that the binomial code does not provide a computational advantage as a starting point for the search.
VI Discussion
An important question when using numerical methods to optimize the operation of quantum systems is how the algorithm scales for large systems. Our algorithm performs a simulation of the time evolution of the open quantum system. This step will follow the scaling laws of simulating a quantum system on a classical computer. In general, it will suffer form the exponential scaling of required resources if we try to apply the algorithm to a many-qudit system. As a result, we can say that our algorithm is most suitable for single- or few qudit systems. This limitation is not a major obstacle to the applicability of the algorithm. As in the case of qubit-based error correction, if the errors on different qudits are uncorrelated, the multi-qubit error rate can be negligibly small, such that it can be ignored when developing the error correction strategy.
It is also worth mentioning that there are methods to extend the range of applicability of classical algorithms to quantum control problems. In particular, machine learning techniques can help speed up the AQEC protocol discovery process Foesel . Another idea that is being discussed in similar contexts is the possible use of hybrid classical-quantum algorithms, where the quantum evolution is determined using a quantum processor while the optimization task is handled by a classical processor Bausch . Our algorithm is well-suited for this approach. Yet another possible way to improve scaling is the use of probabilistic search methods. We recently showed that a probabilistic random search can dramatically enhance the ability of optimal-control methods to find quantum gate decompositions of few-qubit operations Ashhab2024 . Similar ideas could be tried in the AQEC code search problem.